OpenStudy (anonymous):

A ball is kicked upward with an initial velocity of 56 feet per second. The ball's height h (in feet) can be expressed as a function of time t (in seconds) by the equation h = -16t2 + 56t. How much time does the ball take to return to the ground?

OpenStudy (anonymous):

ok, on earth the gravity would be 9.81 m/s^2. Therefore you have the initial velocity. We can call that Vo = 56 ft/s . since the equation is x = Vo*t + 1/2 g*t^2 we can solve it replacing. x = h , -16t^2 + 56t = 56t + 1/2 * -9.81 * t^2. Do the math and solve the cuadratic and there you go..